OpenStudy (anonymous):

Using the properties of equality to solve the equation -2b + 7 = -13, you would _____.

OpenStudy (thedj4jc):

You've just got to do the problem backwards to find the find b... -13 - 7 = ?

OpenStudy (texaschic101):

-2b + 7 = -13 --- subtract 7 from both sides -- (subtraction property) -2b = -13 - 7 -2b = -20 -- divide both sides by -2 --(division property) b = -20/-2 b = 10